# On algebraic properties of some non-prime ideals of collections of cells

### v2

@@ 3. Zig-zag collections
Figure 7. The collection of cells in the right is not a zig-zag collection but contains a zig-zag walk.
@@
Moreover, $I_{\mathcal{P}}$ is not prime because $I_{1},\dots,I_{\ell}$ is a zig-zag walk of $\mathcal{P}$. For more details we refer to [25, Section 3], where the arguments given for polyominoes can be extended easily to collections of cells.
Our first step is to study the Gröbner basis of the inner 2-minors ideal of a zig-zag collection. For this aim, we introduce four total orders on $\mathbb{Z}^{2}$. Let $(i,j),(k,l)\in\mathbb{Z}^{2}$, then we say that:
@@
Let $\mathcal{P}$ be a collection of cells. For $i\in\{1,2,3,4\}$, denote by $<_{\mathrm{lex}}^{(i)}$ the lexicographic order on $K[x_{v}\mid v\in V(\mathcal{P})]$, induced by the following total order on the variables: $x_{a}<_{\mathrm{lex}}^{(i)}x_{b}$ if and only if $a<^{(i)}b$ for $a,b\in V(\mathcal{P})$. If $f$ is an inner 2-minor in $I_{\mathcal{P}}$, observe that for $i\in\{1,3\}$ the initial monomial of $f$ with respect to $<_{\mathrm{lex}}^{(i)}$ is related to the anti-diagonal corners of the inner interval associated. While for $i\in\{2,4\}$ the initial monomial of $f$ is related to the diagonal corners. Moreover, [28, Theorem 4.1] and the arguments of its proof hold also considering the monomial orders $<_{\mathrm{lex}}^{(i)}$ for $i\in\{2,4\}$. The same can be considered for [28, Remark 4.2] with respect to the monomial orders $<_{\mathrm{lex}}^{(i)}$ for $i\in\{1,3\}$. Similar considerations can be made using the reverse lexicographic order, with reference to [26, Proposition 3.2]. For our aim, we state the result below, which can be proved by the same arguments of the mentioned results in [28] or also using the lemmas in [6, Section 3].
@@
In the first step we give a total order on $(I_{1}\setminus\{v_{1}\})\cup I_{2}$. We use $<^{(3)}$ and $<^{(2)}$ respectively to order the vertices in $I_{1}\setminus\{v_{1}\}$ and $I_{2}\setminus\{v_{2}\}$; look at Figure 9b, where the arrows denote the sense of the orders of the vertices in $I_{1}\setminus\{v_{1}\}$ and $I_{2}\setminus\{v_{2}\}$ respectively. Moreover, we set that every vertex of $\mathcal{P}$ in $I_{1}\setminus\{v_{1}\}$ is smaller than each one in $I_{2}\setminus\{v_{2}\}$. Let $k\in[\ell/2]$ with $k\geq 2$ and consider the intervals $I_{2k-2}$, $I_{2k-1}$ and $I_{2k}$. Assume that a total order in $I_{2k-2}\setminus\{v_{2k-2}\}$ is already defined; we want to define a total order on $(I_{2k-1}\setminus\{v_{2k-1}\})\cup I_{2k}$. First of all, we introduce two total orders on $I_{2k-1}\setminus\{v_{2k-1}\}$ and on $I_{2k}\setminus\{v_{2k}\}$ in the following way.
@@
    If $I_{2k-2}$, $I_{2k-1}$ and $I_{2k}$ are arranged as in Figures 10a and 10d or as in Figures 11a, 11b and 11d, then we use $<^{(1)}$ and $<^{(2)}$ to order the vertices in $I_{2k-1}\setminus\{v_{2k-1}\}$ and $I_{2k}\setminus\{v_{2k}\}$ respectively.
@@
    In the case that $I_{2k-2}$, $I_{2k-1}$ and $I_{2k}$ are as in Figure 10b, then we define an order of the vertices in $I_{2k-1}\setminus\{v_{2k-1}\}$ and $I_{2k}\setminus\{v_{2k}\}$ using $<^{(3)}$ and $<^{(2)}$ respectively.
@@
    When $I_{2k-2}$, $I_{2k-1}$ and $I_{2k}$ are in the position of Figure 10c, then use respectively $<^{(1)}$ and $<^{(4)}$ to make order in $I_{2k-1}\setminus\{v_{2k-1}\}$ and $I_{2k}\setminus\{v_{2k}\}$.
@@
We start proving (1). Let $f=x_{a}x_{b}-x_{c}x_{d}$ and $g=x_{p}x_{q}-x_{r}x_{s}$ be two generators of $I_{\mathcal{P}}$ attached to the inner intervals $[a,b]$ and $[p,q]$ of $\mathcal{P}$, respectively. We want to show that the $S$-polynomial $S(f,g)$ of $f$ and $g$ reduces to 00 with respect to $G(\mathcal{P})$. Let $k\in[\ell/2]$. We may consider just the case when $I_{2k-2}$, $I_{2k-1}$ and $I_{2k}$ are arranged as in Figures 10b (with the convention that $I_{0}=I_{\ell}$); indeed, all the other situations described in Figures 10 and 11 can be proved similarly. If $[a,b],[p,q]\subseteq I_{2k-1}$ or $[a,b],[p,q]\subseteq I_{2k}$ then $S(f,g)$ reduces to 00 with respect to $G(\mathcal{P})$ by Proposition 3.3. Assume that $[a,b]\subseteq I_{2k-1}$ and $[p,q]\subseteq I_{2k}$. In such a case, observe that $[a,b]\cap[p,q]=\emptyset$ or $[a,b]\cap[p,q]=\{b\}$ with $b=p$. Since in this case we have also $\operatorname{in}_{<}(f)=x_{c}x_{d}$ and $\operatorname{in}_{<}(g)=x_{p}x_{q}$, then in both cases we obtain $\gcd(\operatorname{in}_{<}(f),\operatorname{in}_{<}(g))=1$. Hence $S(f,g)$ reduces to 00 with respect to $G(\mathcal{P})$. It is trivial to see that $\gcd(\operatorname{in}_{<}(f),\operatorname{in}_{<}(g))=1$ occurs when $[a,b]\subseteq I_{i}$ and $[p,q]\subseteq I_{j}$ with $i,j\in[\ell]$ and $j>i+1.$ In conclusion, the claim (1) is completely proved.
@@
Finally, it is known that if the initial ideal of a homogeneous ideal $I$ of $K[x_{1},\dots,x_{n}]$, with respect to a monomial order, is squarefree then $I$ is radical (see [17, Corollary 2.2]). Since $\mathop{\rm in}\nolimits_{<_{\mathrm{lex}}}(I_{\mathcal{P}})$ is squarefree, then $I_{\mathcal{P}}$ is radical. ∎
@@
Let $\mathcal{P}$ be a zig-zag collection and $\prec$ be the monomial order satisfying the conditions of Lemma 3.4. Denote $\mathcal{E}_{i}=I_{i}\setminus\{v_{i}\}$ for all $i\in[\ell]$. We highlight that, by the proof of Lemma 3.4, the order $\prec$ is the lexicographic order induced by a total order on $V(\mathcal{P})$ and it satisfies the following property: if $a\in\mathcal{E}_{i}$ and $b\in\mathcal{E}_{j}$ with $i<j$ then $x_{a}\prec x_{b}$. This fact is useful further.
For shortness we will write $\mathrm{in}(I_{\mathcal{P}})$ and $\mathrm{in}(I_{\mathcal{P}_{i}})$ for $\mathrm{in}_{\prec}(I_{\mathcal{P}})$ and $\mathrm{in}_{<_{i}}(I_{\mathcal{P}_{i}})$ respectively. Moreover, following the construction in Lemma 3.4, without loss of generality we can always assume that $v_{1}$ is an anti-diagonal corner in $I_{1}$. For all $i\in\{1\ldots,\ell\}$ denote by $S^{\prime}_{i}=K[x_{a}\mid a\in V(I_{i})\setminus\{v_{i}\}]$. Set also $\mathrm{HP}_{K[\mathcal{P}_{i}]}(t)=\frac{h_{i}(t)}{(1-t)^{n_{i}}}$, in particular we know that $n_{i}=|V(\mathcal{P}_{i})|-|\mathcal{P}_{i}|$.
@@
(2) By [28, Theorem 2.2] we know that $K[\mathcal{P}_{i}]$ is a Cohen-Macaulay domain of dimension $|V(\mathcal{P}_{i})|-|\mathcal{P}_{i}|$, so by Proposition 1.6 also $S_{\mathcal{P}_{i}}/\mathrm{in}(I_{\mathcal{P}_{i}})$ is Cohen-Macaulay of dimension $|V(\mathcal{P}_{i})|-|\mathcal{P}_{i}|$. Therefore, by Proposition 1.9 we obtain that $S^{\prime}_{\mathcal{P}_{i}}/\mathrm{in}(I_{\mathcal{P}_{i}})$ is Cohen-Macaulay of dimension $|V(\mathcal{P}_{i})|-|\mathcal{P}_{i}|-1$ and as consequence $S_{\mathcal{P}}/\mathrm{in}(I_{\mathcal{P}})$ is Cohen-Macaulay of dimension $\sum_{i=1}^{\ell}(|V(\mathcal{P}_{i})|-|\mathcal{P}_{i}|-1)=|V(\mathcal{P})|-|\mathcal{P}|$. By Proposition 1.6 the same property holds for $K[\mathcal{P}]$. ∎
Example 3.7.
We provide here an example of a non zig-zag collection of cells, whose coordinate ring is Cohen-Macaulay. Let $\mathcal{P}$ be the collection of cells in Figure 12.
Figure 12. A not zig-zag collection of cells
Denote by $<$ the lexicographic order on $S_{\mathcal{P}}$ induced by the following total order of the variables:
$x_{34}>x_{42}>x_{45}>x_{44}>x_{54}>x_{73}>x_{61}>x_{65}>x_{57}>x_{47}>x_{37}>x_{16}>x_{15}>x_{14}>x_{22}>x_{72}>$
$>x_{66}>x_{64}>x_{63}>x_{62}>x_{56}>x_{55}>x_{53}>x_{52}>x_{51}>x_{46}>x_{43}>x_{36}>x_{33}>x_{32}>x_{27}>$
$>x_{26}>x_{25}>x_{24}>x_{23}>x_{13}$
By some computations we get that the set of the generators of $I_{\mathcal{P}}$ forms the reduced Gröbner basis of $I_{\mathcal{P}}$ with respect to $<$ and the initial ideal of $I_{\mathcal{P}}$ is generated by the monomials:
$x_{42}x_{54},x_{45}x_{54},x_{34}x_{63},x_{44}x_{63},x_{54}x_{63},x_{73}x_{62},x_{65}x_{56},x_{42}x_{55},x_{34}x_{53},x_{42}x_{53},x_{45}x_{53},x_{44}x_{53},x_{61}x_{52},x_{57}x_{46},$
$x_{34}x_{43},x_{57}x_{36},x_{47}x_{36},x_{22}x_{33},x_{57}x_{26},x_{47}x_{26},x_{37}x_{26},x_{16}x_{25},x_{16}x_{24},x_{15}x_{24},x_{16}x_{23},x_{15}x_{23},x_{14}x_{23}.$
Hence $I_{\mathcal{P}}$ is radical but it is not prime since $[(2,2),(3,3)],[(3,3),(6,4)],[(6,2),(7,3)],[(5,1),(6,2)],$ $[(4,2),(5,5)],[(5,5),(6,6)],[(2,6),(5,7)],[(1,3),(2,6)]$ is a zig-zag walk of $\mathcal{P}$. Using Macaulay2 ([4, 12]) we check that $S_{\mathcal{P}}/\mathrm{in}_{<}(I_{\mathcal{P}})$ is a Cohen-Macaulay ring. Hence, from Proposition 1.6, we have that $K[\mathcal{P}]$ is Cohen-Macaulay.
We point out that we needed to check the Cohen-Macaulayness of $S_{\mathcal{P}}/\mathrm{in}_{<}(I_{\mathcal{P}})$ and not of $S_{\mathcal{P}}/I_{\mathcal{P}}$ using Macaulay2, because the time of computation to check the same property for $S_{\mathcal{P}}/I_{\mathcal{P}}$ is too long.
@@
For $i\in[\ell]$ we denote by $h_{i}(t)$ the $h$-polynomial of $K[\mathcal{P}_{i}]$, so from [29, Theorem 3.5] we have that $h_{i}(t)$ is the switching rook-polynomial of $\mathcal{P}_{i}$ for all $i\in[\ell]$. Recall that Theorem 3.6 states in particular that $h(t)=\prod_{i=1}^{\ell}h_{i}(t)$. From Lemma 1.1 we have that $\prod_{i=1}^{\ell}h_{i}(t)$ is the switching rook polynomial of $\cup_{i=1}^{\ell}\mathcal{P}_{i}$, that is $h(t)$ is the switching rook polynomial of $\mathcal{P}$. Moreover, since $K[\mathcal{P}]$ is Cohen-Macaulay from Theorem 3.6, then it follows that $\mathrm{reg}(K[\mathcal{P}])=r(\mathcal{P})$. ∎
@@
(1) Let $h(t)=\sum_{k=1}^{s}h_{k}t^{k}$ be the $h$-polynomial of $K[\mathcal{P}]$. From Corollary 3.8 we have that $h(t)$ is the switching rook polynomial of $\mathcal{P}$ and $s=r(\mathcal{P})$. Since $K[\mathcal{P}]$ is Gorenstein, it follows from [37, Corollary 5.3.10] that $h_{k}=h_{s-k}$ for all $k\in[s]$. In particular, we have that $h_{s}=1$. Suppose by contradiction that there exists $h\in[\ell]$ such that $\mathcal{P}_{h}$ is not a square. Up to reflections, rotations or translations of $\mathcal{P}$, we may assume that $\mathcal{P}_{h}=\mathcal{P}_{[(1,1),(m.n)]}$, where $m>n$. Observe that $r(\mathcal{P})=n-1$. We define a canonical configuration $\mathcal{T}_{1}$ of $n-1$ rooks in $\mathcal{P}_{h}$, placing a rook in the cell of $\mathcal{P}_{h}$ having lower left corner $(j,j)$ for all $j\in[n-1]$. Moreover, since $m>n$, we can define another canonical configuration $\mathcal{T}_{2}$ of $n-1$ rooks in $\mathcal{P}_{h}$, placing a rook in the cell of $\mathcal{P}_{h}$ having lower left corner $(j,j)$ for all $j\in[n-2]$ and a rook in that one with $(n,n-1)$ as lower left corner. Denote by $\mathcal{C}_{i}$ a canonical configuration of $r(\mathcal{P}_{i})$ rooks in $\mathcal{P}_{i}$, for all $i\in[s]\setminus\{h\}$, and set $\mathcal{C}=\cup_{i\in[s]\setminus\{h\}}\mathcal{C}_{i}$. It is easy to see that $\mathcal{C}\cup\mathcal{T}_{1}$ and $\mathcal{C}\cup\mathcal{T}_{2}$ are two canonical configuration of $r(\mathcal{P})$ rooks in $\mathcal{P}$, so $h_{s}>1$. This is a contradiction because $h_{s}=1$. In conclusion $\mathcal{P}_{i}$ is a square for all $i\in[\ell]$.
(2) Observe that $S_{\mathcal{P}}/\mathrm{in}(I_{\mathcal{P}})$ can be viewed as an edge ring of a graph $G$ on the vertex set $V(\mathcal{P})$, where $\{a,b\}\in E(G)$ if either $a,b$ are the anti-diagonal corners of an inner interval of $I_{2k}$ or they are the diagonal corners of an inner interval of $I_{2k-1}$, for $k\in[\ell/2]$. If $\mathcal{P}_{i}$ is a cell for all $i\in[\ell]$, $G$ is a disjoint union of edges, so $S_{\mathcal{P}}/\mathrm{in}(I_{\mathcal{P}})$ is Gorenstein by [20, Corollary 9.3.3]. Therefore, it follows that $K[\mathcal{P}]$ is Gorenstein by (2) of Proposition 1.5. ∎
@@
If there exists $j\in[\ell]$ such that $\mathcal{P}_{j}$ is a square different from a cell, then $S_{\mathcal{P}}/\mathrm{in}(I_{\mathcal{P}})$ cannot be Gorenstein because $G$ is a chordal graph but it is not a disjoint union of edges. Anyway, we tested different examples with Macaulay2 that shows $K[\mathcal{P}]$ is still Gorenstein. So, we ask if the converse of (1) is true in general.
4. Cohen-Macaulay property and Hilbert series for closed path polyominoes having zig-zag walks
From now onward, we assume that $\mathcal{P}$ is a non-prime closed path polyomino. By [3, Theorem 6.2], $\mathcal{P}$ contains a zig-zag walk. See Figure 13 for an example of a closed path with zig-zag walks.
Figure 13. Closed path with zig-zag walks.
{tikzpicture}
{tikzpicture}

### v3

@@ 3. Zig-zag collections
Figure 5. Collections of cells containing a zig-zag walk.
@@
Moreover, $I_{\mathcal{P}}$ is not prime because $I_{1},\dots,I_{\ell}$ is a zig-zag walk of $\mathcal{P}$. For more details we refer to [26, Section 3], where the arguments given for polyominoes can be extended easily to collections of cells.
Our first step is to study the Gröbner basis of the inner 2-minor ideal of a zig-zag collection. For this aim, we introduce four total orders on $\mathbb{Z}^{2}$. Let $(i,j),(k,l)\in\mathbb{Z}^{2}$, then we say that:
@@
Let $\mathcal{P}$ be a collection of cells. For $i\in\{1,2,3,4\}$, denote by $<_{\mathrm{lex}}^{(i)}$ the lexicographic order on $K[x_{v}\mid v\in V(\mathcal{P})]$, induced by the following total order on the variables: $x_{a}<_{\mathrm{lex}}^{(i)}x_{b}$ if and only if $a<^{(i)}b$ for $a,b\in V(\mathcal{P})$. If $f$ is an inner 2-minor in $I_{\mathcal{P}}$, observe that for $i\in\{1,3\}$ the initial monomial of $f$ with respect to $<_{\mathrm{lex}}^{(i)}$ is related to the anti-diagonal corners of the associated inner interval. While for $i\in\{2,4\}$ the initial monomial of $f$ is related to the diagonal corners. Moreover, [29, Theorem 4.1] and the arguments of its proof hold also considering the monomial orders $<_{\mathrm{lex}}^{(i)}$ for $i\in\{2,4\}$. The same can be considered for [29, Remark 4.2] with respect to the monomial orders $<_{\mathrm{lex}}^{(i)}$ for $i\in\{1,3\}$. Similar considerations can be made using the reverse lexicographic order, with reference to [27, Proposition 3.2]. For our aim, we state the result below, which can be proved by the same arguments of the mentioned results in [29] or also using the lemmas in [7, Section 3].
@@
In the first step we give a total order on $(I_{1}\setminus\{v_{1}\})\cup I_{2}$. We use $<^{(3)}$ and $<^{(2)}$, respectively, to order the vertices in $I_{1}\setminus\{v_{1}\}$ and $I_{2}\setminus\{v_{2}\}$; look at Figure 7b, where the arrows denote the sense of the orders of the vertices in $I_{1}\setminus\{v_{1}\}$ and $I_{2}\setminus\{v_{2}\}$ respectively. Moreover, we set that every vertex of $\mathcal{P}$ in $I_{1}\setminus\{v_{1}\}$ is smaller than each one in $I_{2}\setminus\{v_{2}\}$. Let $k\in[\ell/2]$ with $k\geq 2$ and consider the intervals $I_{2k-2}$, $I_{2k-1}$ and $I_{2k}$. Assume that a total order in $I_{2k-2}\setminus\{v_{2k-2}\}$ is already defined; we want to define a total order on $(I_{2k-1}\setminus\{v_{2k-1}\})\cup I_{2k}$. First of all, we introduce two total orders on $I_{2k-1}\setminus\{v_{2k-1}\}$ and on $I_{2k}\setminus\{v_{2k}\}$ in the following way.
@@
    If $I_{2k-2}$, $I_{2k-1}$ and $I_{2k}$ are arranged as in Figures 8a and 8d or as in Figures 9a, 9b and 9d, then we use $<^{(1)}$ and $<^{(2)}$ to order the vertices in $I_{2k-1}\setminus\{v_{2k-1}\}$ and $I_{2k}\setminus\{v_{2k}\}$ respectively.
@@
    In the case that $I_{2k-2}$, $I_{2k-1}$ and $I_{2k}$ are as in Figure 8b, then we define an order of the vertices in $I_{2k-1}\setminus\{v_{2k-1}\}$ and $I_{2k}\setminus\{v_{2k}\}$ using $<^{(3)}$ and $<^{(2)}$, respectively.
@@
    When $I_{2k-2}$, $I_{2k-1}$ and $I_{2k}$ are in the position of Figure 8c, then use respectively $<^{(1)}$ and $<^{(4)}$ to make an order in $I_{2k-1}\setminus\{v_{2k-1}\}$ and $I_{2k}\setminus\{v_{2k}\}$.
@@
We start by proving (1). Let $f=x_{a}x_{b}-x_{c}x_{d}$ and $g=x_{p}x_{q}-x_{r}x_{s}$ be two generators of $I_{\mathcal{P}}$ attached to the inner intervals $[a,b]$ and $[p,q]$ of $\mathcal{P}$, respectively. We want to show that the $S$-polynomial $S(f,g)$ of $f$ and $g$ reduces to 00 with respect to $G(\mathcal{P})$. Let $k\in[\ell/2]$. We may consider just the case when $I_{2k-2}$, $I_{2k-1}$ and $I_{2k}$ are arranged as in Figures 8b (with the convention that $I_{0}=I_{\ell}$); indeed, all the other situations described in Figures 8 and 9 can be proved similarly. If $[a,b],[p,q]\subseteq I_{2k-1}$ or $[a,b],[p,q]\subseteq I_{2k}$ then $S(f,g)$ reduces to 00 with respect to $G(\mathcal{P})$ by Proposition 3.3. Assume that $[a,b]\subseteq I_{2k-1}$ and $[p,q]\subseteq I_{2k}$. In such a case, observe that $[a,b]\cap[p,q]=\emptyset$ or $[a,b]\cap[p,q]=\{b\}$ with $b=p$. Since in this case we have also $\operatorname{in}_{<}(f)=x_{c}x_{d}$ and $\operatorname{in}_{<}(g)=x_{p}x_{q}$, then in both cases we obtain $\gcd(\operatorname{in}_{<}(f),\operatorname{in}_{<}(g))=1$. Hence $S(f,g)$ reduces to 00 with respect to $G(\mathcal{P})$. It is trivial to see that $\gcd(\operatorname{in}_{<}(f),\operatorname{in}_{<}(g))=1$ occurs when $[a,b]\subseteq I_{i}$ and $[p,q]\subseteq I_{j}$ with $i,j\in[\ell]$ and $j>i+1.$ In conclusion, the claim (1) is completely proved.
@@
Finally, it is known that if the initial ideal of a homogeneous ideal $I$ of $K[x_{1},\dots,x_{n}]$, with respect to a monomial order, is squarefree then $I$ is radical (see [18, Corollary 2.2]). Since $\mathop{\rm in}\nolimits_{<_{\mathrm{lex}}}(I_{\mathcal{P}})$ is squarefree, then $I_{\mathcal{P}}$ is radical. ∎
@@
Let $\mathcal{P}$ be a zig-zag collection and $\prec$ be the monomial order satisfying the conditions of Lemma 3.4. Denote $\mathcal{E}_{i}=I_{i}\setminus\{v_{i}\}$ for all $i\in[\ell]$. We highlight that, by the proof of Lemma 3.4, the order $\prec$ is the lexicographic order induced by a total order on $V(\mathcal{P})$ and it satisfies the following property: if $a\in\mathcal{E}_{i}$ and $b\in\mathcal{E}_{j}$ with $i<j$ then $x_{a}\prec x_{b}$. This observation will be useful in later arguments.
For shortness, we will write $\mathrm{in}(I_{\mathcal{P}})$ and $\mathrm{in}(I_{\mathcal{P}_{i}})$ for $\mathrm{in}_{\prec}(I_{\mathcal{P}})$ and $\mathrm{in}_{<_{i}}(I_{\mathcal{P}_{i}})$ respectively. Moreover, following the construction in Lemma 3.4, without loss of generality we can always assume that $v_{1}$ is an anti-diagonal corner in $I_{1}$. For all $i\in\{1\ldots,\ell\}$ denote by $S^{\prime}_{i}=K[x_{a}\mid a\in V(I_{i})\setminus\{v_{i}\}]$. Set also $\mathrm{HP}_{K[\mathcal{P}_{i}]}(t)=\frac{h_{i}(t)}{(1-t)^{n_{i}}}$, in particular we know that $n_{i}=|V(\mathcal{P}_{i})|-|\mathcal{P}_{i}|$. Finally, before proving the main result of this section, we recall a beautiful result achieved in [10], which will be crucial also in the other parts of this work.
Proposition 3.6 ([10]).
Let $I$ be a homogeneous ideal in a polynomial ring $S$ and suppose that $\mathrm{in}(I)$ is radical with respect to some term order. Then $\mathrm{reg}(S/I)=\mathrm{reg}(S/\mathrm{in}(I))$ and $\mathrm{depth}(S/I)=\mathrm{depth}(S/\mathrm{in}(I))$. Furthermore, $S/I$ is Cohen-Macaulay if and only if $S/\mathrm{in}(I)$ is Cohen-Macaulay.
@@
(2) By [29, Theorem 2.2] we know that $K[\mathcal{P}_{i}]$ is a Cohen-Macaulay domain of dimension $|V(\mathcal{P}_{i})|-|\mathcal{P}_{i}|$, so by Proposition 3.6 also $S_{\mathcal{P}_{i}}/\mathrm{in}(I_{\mathcal{P}_{i}})$ is Cohen-Macaulay of dimension $|V(\mathcal{P}_{i})|-|\mathcal{P}_{i}|$. Therefore, by [37, Lemma 3.1.35] we obtain that $S^{\prime}_{\mathcal{P}_{i}}/\mathrm{in}(I_{\mathcal{P}_{i}})$ is Cohen-Macaulay of dimension $|V(\mathcal{P}_{i})|-|\mathcal{P}_{i}|-1$ and, as a consequence, $S_{\mathcal{P}}/\mathrm{in}(I_{\mathcal{P}})$ is Cohen-Macaulay of dimension $\sum_{i=1}^{\ell}(|V(\mathcal{P}_{i})|-|\mathcal{P}_{i}|-1)=|V(\mathcal{P})|-|\mathcal{P}|$. By Proposition 3.6 the same property holds for $K[\mathcal{P}]$. ∎
@@
For $i\in[\ell]$ we denote by $h_{i}(t)$ the $h$-polynomial of $K[\mathcal{P}_{i}]$, so from [30, Theorem 3.5] we have that $h_{i}(t)$ is the switching rook-polynomial of $\mathcal{P}_{i}$ for all $i\in[\ell]$. Recall that Theorem 3.7 states in particular that $h(t)=\prod_{i=1}^{\ell}h_{i}(t)$. From Lemma 1.1 we have that $\prod_{i=1}^{\ell}h_{i}(t)$ is the switching rook polynomial of $\cup_{i=1}^{\ell}\mathcal{P}_{i}$, that is $h(t)$ is the switching rook polynomial of $\mathcal{P}$. Moreover, since $K[\mathcal{P}]$ is Cohen-Macaulay from Theorem 3.7, then it follows that $\mathrm{reg}(K[\mathcal{P}])=r(\mathcal{P})$. ∎
@@
(1) Let $h(t)=\sum_{k=1}^{s}h_{k}t^{k}$ be the $h$-polynomial of $K[\mathcal{P}]$. From Corollary 3.8 we have that $h(t)$ is the switching rook polynomial of $\mathcal{P}$ and $s=r(\mathcal{P})$. Since $K[\mathcal{P}]$ is Gorenstein, it follows from [37, Corollary 5.3.10] that $h_{k}=h_{s-k}$ for all $k\in[s]$. In particular, we have that $h_{s}=1$. Suppose by contradiction that there exists $h\in[\ell]$ such that $\mathcal{P}_{h}$ is not a square. Up to reflections, rotations or translations of $\mathcal{P}$, we may assume that $\mathcal{P}_{h}=\mathcal{P}_{[(1,1),(m.n)]}$, where $m>n$. Observe that $r(\mathcal{P})=n-1$. We define a canonical configuration $\mathcal{T}_{1}$ of $n-1$ rooks in $\mathcal{P}_{h}$, placing a rook in the cell of $\mathcal{P}_{h}$ having lower left corner $(j,j)$ for all $j\in[n-1]$. Moreover, since $m>n$, we can define another canonical configuration $\mathcal{T}_{2}$ of $n-1$ rooks in $\mathcal{P}_{h}$, placing a rook in the cell of $\mathcal{P}_{h}$ having lower left corner $(j,j)$ for all $j\in[n-2]$ and a rook in that one with $(n,n-1)$ as lower left corner. Denote by $\mathcal{C}_{i}$ a canonical configuration of $r(\mathcal{P}_{i})$ rooks in $\mathcal{P}_{i}$, for all $i\in[s]\setminus\{h\}$, and set $\mathcal{C}=\cup_{i\in[s]\setminus\{h\}}\mathcal{C}_{i}$. It is easy to see that $\mathcal{C}\cup\mathcal{T}_{1}$ and $\mathcal{C}\cup\mathcal{T}_{2}$ are two canonical configurations of $r(\mathcal{P})$ rooks in $\mathcal{P}$, so $h_{s}>1$. This is a contradiction because $h_{s}=1$. In conclusion $\mathcal{P}_{i}$ is a square for all $i\in[\ell]$.
(2) Observe that $S_{\mathcal{P}}/\mathrm{in}(I_{\mathcal{P}})$ can be viewed as an edge ring of a graph $G$ on the vertex set $V(\mathcal{P})$, where $\{a,b\}\in E(G)$ if either $a,b$ are the anti-diagonal corners of an inner interval of $I_{2k}$ or they are the diagonal corners of an inner interval of $I_{2k-1}$, for $k\in[\ell/2]$. If $\mathcal{P}_{i}$ is a cell for all $i\in[\ell]$, $G$ is a disjoint union of edges, so $S_{\mathcal{P}}/\mathrm{in}(I_{\mathcal{P}})$ is Gorenstein by [21, Corollary 9.3.3]. Therefore, it follows that $K[\mathcal{P}]$ is Gorenstein by [37, Proposition 9.6.17]. ∎
@@
If there exists $j\in[\ell]$ such that $\mathcal{P}_{j}$ is a square different from a cell, then $S_{\mathcal{P}}/\mathrm{in}(I_{\mathcal{P}})$ cannot be Gorenstein because $G$ is a chordal graph but it is not a disjoint union of edges. Anyway, we tested different examples with Macaulay2 ([5, 13]) that shows $K[\mathcal{P}]$ is still Gorenstein. So, we ask if the converse of (1) is true in general.
4. Cohen-Macaulay property and Hilbert-Poincaré series of closed paths with zig-zag walks
In this section we investigate Cohen-Macaulayness and the Hilbert-Poincaré series of non-prime closed path polyominoes. From now onward, we assume that $\mathcal{P}$ is a non-prime closed path. By [4, Theorem 6.2], $\mathcal{P}$ contains a zig-zag walk. See Figure 3b for an example of a closed path with zig-zag walks.
Discussion 4.1.
If $\mathcal{P}$ is a closed path polyomino having zig-zag walks, by [4, Proposition 6.1] $\mathcal{P}$ does not contain any $L$-configuration or ladders with at least three steps. So, any two maximal inner intervals of $\mathcal{P}$ intersect themselves in the cells displayed in Figure 10a or 10b (up to reflections or rotations).
$m$$X$
$W$$X$$Y$$Z$